# Scale-aware Recognition in Satellite Images under Resource Constraints

### v1

@@ 4.2.3 Model Disagreement for Modality Selection
Following the methodology in sec. 3.5, we describe how the disagreement between HR and LR models can act as a good indicator for which locations would benefit from HR imagery. We demonstrate how well the KD model performs as a stand in for the HR model, so that we may calculate this disagreement without requiring HR imagery. Figure 3 visualizes images ranked by degree of disagreement between the LR and HR model (top) and between the LR and KD model (bottom). We see that both approaches yield similar rankings. Therefore the disagreement between the LR and KD model is a good proxy for locations where HR data is needed.
Figure 3: Images ranked according to disagreement between the LR and HR model (top) and the LR and KD model (bottom). Both rankings are similar, even though the latter only uses LR images.
@@
We compare the performance of our sampling technique against the sampling technique presented in [27], in which night lights are used as an indicator to sample more densely. The results of this sampling technique are included in Table 2.
@@ 4.2.4 Concept Scale Inference
We compared our LLM-based approach for inferring the best modality per concept with other baseline approaches. These baselines include either always choosing the HR modality, always choosing the LR modality, or selecting between the two based on the average area of each concept as measured in the Open Street Maps (OSM) annotations. For each approach, we evaluate its accuracy in terms of determining the right modality.
We evaluate this on 10 unseen classes as the seen concepts were given to the LLM as examples as described in sec. 6.2. The LLM approach selected the correct resolution 100% of the time. Both the OSM average area approach and choosing only HR imagery were correct  90% of the time whereas selecting solely LR imagery was correct only  10% of the time. This result is interesting as the LLM outperformed using the average area of the concepts (OSM), suggesting that it is not just the size of the concept that is an important consideration, but also a semantic understanding of the concept and its features. We include a tabulated version of these results in the appendix, sec. 5.
@@
We also utilize OSM data to train and evaluate our models. We acknowledge that this data contains noise, due to its crowd-sourced nature. Unfortunately, there are not many sources for fine-grained labeling of satellite imagery. Therefore, we follow past work in using OSM for training and evaluation [26, 3]. Moreover other standard sources of labeled satellite imagery such as Microsoft’s building footprint dataset work with similar noise levels [4]. Additionally, to show robustness to this noise, we performed an experiment using the GRAFT LR model, wherein we constructed bootstrap samples by sub-sampling within our original test set 1000 times. The results show that the 90% confidence interval is within 2.43% of the reported mAP.

### v2

@@ 4.2.3 Model Disagreement for Modality Selection
Following the methodology in sec. 3.5, we describe how the disagreement between HR and LR models can act as a good indicator for which locations would benefit from HR imagery. We found that correlation coefficient between the disagreement scores of the HR (NAIP) vs KD with the LR model is 0.9322, which signifies a strong positive correlation. Similarly, the correlation coefficient between the disagreement scores of the HR (NICFI) vs KD with the LR model is 0.998, which also signifies a strong positive correlation.
Figure 3 visualizes images ranked by degree of disagreement between the LR and HR model (top) and between the LR and KD model (bottom). We see that both approaches yield similar rankings demonstrating that the KD model can act a stand in for the HR model, so that we may calculate the disagreement without requiring HR imagery.
Figure 3: Images ranked according to disagreement between the LR and HR model (top) and the LR and KD model (bottom). Both rankings are similar, with a correlation coefficient of 0.9322, even though the latter only uses LR images.
@@
We compare the performance of our sampling technique against the sampling technique presented in (Meng et al., 2022), in which night lights are used as an indicator to sample more densely. The results of this sampling technique are included in Table 2.
@@ 4.2.4 Concept Scale Inference
We compare our LLM-based approach for inferring the best modality per concept with other baseline approaches. For each approach, we evaluate its accuracy in terms of determining the correct modality. These baselines include either always choosing the HR modality and always choosing LR modality, or selecting between the two based on the average area of each concept, as provided via OpenStreetMaps (OSM). Our LLM approach (tested using several off-the-shelf models) achieves 100% accuracy when determining the correct modality our set of concepts. We test on both seen and unseen, demonstrating the ability of easily being extended to far more concepts.
Both the OSM average area approach and choosing only HR imagery were correct  90% of the time whereas selecting solely LR imagery was correct only  10% of the time. This result is interesting as the LLM outperformed using the average area of the concepts (OSM), suggesting that it is not just the size of the concept that is an important consideration, but also a semantic understanding of the concept and its features. We include a tabulated version of these results in the appendix  A.3.
Additional Experiments. We include the following additional experiments in the Appendices to further exhibit the effectiveness of our approach. We demonstrate the impact of different budgets, ranging from 100 locations to 1000 in A.4. Additionally, we compare our KD model with the HR model on Super-Resolution images to calculate model disagreement scores in A.5.
@@
We also utilize OSM data to train and evaluate our models. We acknowledge that this data contains noise, due to its crowd-sourced nature. Unfortunately, there are not many sources for fine-grained labeling of satellite imagery. Therefore, we follow past work in using OSM for training and evaluation (Mall et al., 2024; Bastani et al., 2023). Moreover other standard sources of labeled satellite imagery such as Microsoft’s building footprint dataset work with similar noise levels (Bing Maps, 2018). Additionally, to show robustness to this noise, we performed an experiment using the GRAFT LR model, wherein we constructed bootstrap samples by sub-sampling within our original test set 1000 times. The results show that the 90% confidence interval is within 2.43% of the reported mAP.
@@ 5 Conclusion
Appendix A Appendix
A.1 Multi-Label Classes
The classes we use for labeling our data as well as for testing are the following.
’tennis’, ’skate’, ’american football’, ’swimming’, ’cemetery’, ’garage’, ’golf’, ’roundabout’, ’parking lot’, ’supermarket’, ’school’, ’marina’, ’baseball’, ’fall’, ’pond’, ’airport’, ’beach’, ’bridge’, ’religious’, ’residential’, ’warehouse’, ’office’, ’farmland’, ’university’, ’forest’, ’lake’, ’nature reserve’, ’park’, ’sand’, ’soccer’, ’equestrian’, ’shooting’, ’ice rink’, ’commercial area’, ’garden’, ’dam’, ’railroad’, ’highway’, ’river’, ’wetland’
In the case of the unsupervised techniques, the first 30 are used as the seen classes, while the latter 10 are treated as unseen classes.
A.2 Prompts to LLM used when Inferring Concept Scale
The prompt was “Please act as a binary classifier for the following concepts to determine if they are better suited for ’LR’ (low resolution) or ’HR’ (high resolution) imagery. I will first give you some examples with the correct response. Then given a concept you are to return ’lr’ or ’hr’”
This was followed by the list of ‘seen’ concepts and their optimal resolution in the format “concept:resolution”. After this the LLM only responded with either ‘lr’ or ‘hr’ when given a concept.
A.3 Tabulated Concept Scale Inference Results
Seen Classes Unseen Classes
Technique Accuracy Accuracy
LLM ChatGPT 3.5 (ours) 100% 100%
LLM Gemini (ours) 100% 100%
LL Claude (ours) 100% 100%
OSM 73.33% 90%
Always HR 83.33% 90%
Always LR 16.67% 10%
Table 5: Comparison of concept recognition performance (accuracy, precision, recall) between baseline models and LLM selected models.
We compared our LLM-based approach for inferring the best modality per concept with other baseline approaches. For each approach, we evaluate its accuracy in terms of determining the right modality. These baselines include either always choosing the HR modality and always choosing LR modality, or selecting between the two based on the average area of each concept, as provided via OpenStreetMaps (OSM) OpenStreetMap contributors (2024).
We evaluate on the 10 unseen classes as the seen concepts were given to the LLM as examples as described in sec. A.2.
The LLM approach selected the correct resolution 100% of the time. Both the OSM average area approach and choosing only HR imagery were correct  90%. Finding that the LLM outperformed using the average area of the concepts (OSM), suggests that it is not just the size of the concept that is an important consideration, but also a semantic understanding of the concept itself.
A.4 Impact of Different Budgets
We are able to get an improved accuracy on a tight budget since we only sample HR images when necessary. The reason we are able to do so is (a) some concepts are better in LR and (b) for the remaining concepts we are able to recognize which regions require HR. Of course if the budget decreases further, at some point, the accuracy will also decrease. We demonstrate this in table 6.
Seen Classes* Unseen Classes # HR images Inf. Time
Model Data mAP100 mAP20 mAP100 mAP20 sq. km. s
GRAFT HR HR 0.501 0.513 0.541 0.574 25,163 1539
GRAFT LR LR 0.482 0.507 0.379 0.471 0 17
GRAFT (Ours full), Budget 100 LR,HR 0.557 0.605 0.439 0.521 595 48
GRAFT (Ours full), Budget 500 LR,HR 0.601 0.626 0.480 0.542 2,977 170
GRAFT (Ours full), Budget 750 LR,HR 0.614 0.617 0.495 0.557 4,466 247
GRAFT (Ours full), Budget 1000 LR,HR 0.633 0.639 0.502 0.564 5,954 372
Table 6: Performance of our overall technique utilizing various budgets. *: Classes seen during knowledge distillation, for all other zero-shot baselines all classes are unseen. We are able to get an improved accuracy on a tight budget since we only sample HR images when necessary.
A.5 Use of Super Resolution model for Model Disagreement Score
Here we present the results using the SR images with the HR model to calculate the model disagreement score to sample HR imagery.
Generating and performing inference using the SR data proved to be significantly more computationally expensive. Here we report inference time, one can see the process is significantly slower than using the KD model on the LR images. This is due to the fact that the SR model accepts 32x32 pixel patches of Sentinel-2 images at a time. So for our test set of 5015 LR images of size 256x256 pixels, the SR model is performing inference on 320,960 patches. This process took 12,549.90 s ( 3.5 hours). Additionally upon stitching these SR patches to recreate the coverage of the original LR image, we must split the stitched SR image into 100 components corresponding to the same organization of HR images. This leads to inference using the HR model on 501,500 images. And even after this, one would still need to run the KD model for regions not covered by the budget. Overall this significant jump in inference time makes using any superresolution model in place of our KD model less favorable.
Additionally, our KD model is specifically trained to identify cases where HR data would provide meaningful improvements in recognition accuracy. In contrast, SR models attempt to reconstruct all high-frequency details, regardless of their importance to the recognition task. This targeted approach makes our method more efficient at identifying truly necessary cases for HR image use.
Here we present the results using the SR images with the HR model to calculate the model disagreement score to sample HR imagery. LR + SR + LLM denotes the use of both the LLM and model disagreement components of our work, (using the SR model disagreement), without the use of any KD model. KD + SR + LLM denotes results with our full system, but with the model disagreement scores being obtained via the SR imagery.
While it does offer improvements in mAP for unseen classes, the inference time alone is 34 times longer than our technique, and this improvement only comes with the use of our KD model.
Seen Classes* Unseen Classes # HR images Inf. Time
Model Data mAP100 mAP20 mAP100 mAP20 sq. km. s
GRAFT HR HR 0.501 0.513 0.541 0.574 25,163 1539
GRAFT LR + SR + LLM LR,HR 0.527 0.583 0.512 0.547 5,954 12,867
GRAFT KD + SR + LLM LR,HR 0.554 0.610 0.556 0.589 5,954 12,884
GRAFT (Ours full) LR,HR 0.633 0.639 0.502 0.564 5,954 372
Table 7: Comparison between KD model for disagreement score calculation and HR model on Super Resolution (SR) imagery.:Techniques utilizing our contributions. *: Classes seen during knowledge distillation, for all other zero-shot baselines all classes are unseen. While SR does offer improvements in mAP for unseen classes, the inference time alone is 34 times longer than our technique, and this improvement only comes with the use of our KD model.
A.6 Results Per Concept
Table 8 contains results for each class, allowing one to see which models performed best between the supervised HR model and the novel technique presented in this paper. One can see that for most classes, our model manages to out perform the HR model, without needing nearly as many HR images, and for some concepts with no HR images at all.
Technique
Supervised HR Supervised Ours Full
Concept mAP100 mAP20 # HR imgs mAP100 mAP20 # HR imgs
Tennis 0.617 0.701 25,163 0.969 1.0 5,954
Skate Park 0.277 0.300 25,163 0.163 0.345 5,954
American Football 0.449 0.663 25,163 0.265 0.305 5,954
Swimming 0.789 0.781 25,163 0.912 0.879 5,954
Cemetery 0.294 0.308 25,163 0.615 0.817 0
Garage 0.696 0.835 25,163 0.964 1.0 5,954
Golf 0.483 0.194 25,163 0.974 1.0 0
Roundabout 0.401 0.471 25,163 0.823 0.984 5,954
Parking Lot 1.0 1.0 25,163 1.0 1.0 5,954
Supermarket 0.928 0.978 25,163 0.773 0.756 5,954
School 0.824 0.717 25,163 0.913 0.956 5,954
Marina 0.389 0.516 25,163 0.189 0.294 5,954
Baseball 0.672 0.747 25,163 0.658 0.713 5,954
Pond 0.921 0.967 25,163 0.888 0.923 5,954
Airport 0.558 0.843 25,163 0.590 0.934 5,954
Beach 0.970 0.995 25,163 0.910 0.921 5,954
Bridge 0.986 1.0 25,163 0.976 1.0 5,954
Religious Buildings 0.933 1.0 25,163 0.980 1.0 5,954
Residential 1.0 1.0 25,163 1.0 1.0 5,954
Warehouse 0.901 0.969 25,163 0.945 0.927 5,954
Office 0.955 0.995 25,163 0.961 0.914 0
Farmland 0.859 0.883 25,163 0.854 0.883 0
University 0.505 0.448 25,163 0.652 0.920 0
Forest 0.977 0.995 25,163 0.933 1.0 0
Lake 0.206 0.458 25,163 0.257 0.496 5,954
Nature Reserve 0.938 0.997 25,163 0.915 0.995 5,954
Park 1.0 1.0 25,163 0.983 0.997 0
Sand Pits 0.847 0.811 25,163 0.983 1.0 5,954
Soccer 0.497 0.440 25,163 0.122 0.053 5,954
Equestrian 0.395 0.497 25,163 0.083 0.200 5,954
Shooting Range 0.01 0.05 25,163 0.01 0.05 5,954
Commercial Area 0.483 0.425 25,163 0.895 0.934 5,954
Garden 0.815 0.839 25,163 0.901 0.956 0
Dam 0.062 0.227 25,163 0.062 0.227 5,954
Railroad 0.964 0.974 25,163 0.900 0.974 5,954
Highway 1.0 1.0 25,163 1.0 1.0 0
River 1.0 1.0 25,163 0.967 1.0 5,954
Wetland 0.827 0.925 25,163 0.824 0.957 5,954
Table 8: Performance of various algorithms for various concepts. One can see that our usage of HR images is fully concept aware, and we are able to even outperform models that use solely HR imagery.